OpenStudy (anonymous):

what is the volume of a liquid with a density of 1.5g/ml and a mass of 78g?

OpenStudy (mathstudent55):

Density is mass divided by volume, or d = m/v If you solve for volume, you get v = m/d Just plug in your values for density and mass to find the volume.

OpenStudy (anonymous):

so v+ 78/1.5?

OpenStudy (anonymous):

v= i mean

OpenStudy (anonymous):

so would it be 52?

OpenStudy (mathstudent55):

Yes, but don't forget the units.

OpenStudy (anonymous):

so 52g/ml?

OpenStudy (mathstudent55):

No. Remember you are looking for a volume. Of all units you see in this problem, which units are units of volume? There is no need to guess. If you write the units in the equation, you'll get the correct units in the end.

OpenStudy (anonymous):

g2/ml

OpenStudy (mathstudent55):

\(v = \dfrac{m}{d} \) \(v = \dfrac{78 ~g}{1.5 \frac{g}{ml}} \) \(v = 52 ~g \times \dfrac{ml}{g} \) \(v = 52 ~\cancel{g} \times \dfrac{ml}{\cancel{g}} \) \(v = 52 ~ml\)

OpenStudy (mathstudent55):

ml is a unit of volume g is a unit of mass g/ml is a unit of density

OpenStudy (anonymous):

but u dont add th eunit for volume?